40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 202 of 344  Not logged in ELOG logo
ID Date Author Type Category Subject
  7169   Tue Aug 14 04:32:49 2012 rana, yoichiUpdateLockingPOX signal sometimes looks very funny

 The alignment was way off. We moved the PZT, the BS, and the x arm to get it to lock. Along the way we noticed that giving the ETM and POS offsets makes it tilt a lot. The DC coil balancing is no good at all.

After locking, we tuned up the X arm filters in the LSC and activated the filter module triggers.  I would attach a screenshot of the trigger screen, but sadly it has no snapshot button on it.

WE changed the integrator into a double integrator with a complex zero pair. We also replaced the 1:50 boost with a 2nd order complex pole:zero pair. And added a 18 Hz RG. These were all set by looking at the error point spectra and minimizing the RMS. Hopefully, this kind of work will all be obsolete once we get the optimal feedback code. For now, the arm is very stable - we're leaving it locked overnight since the filter triggering seems to work well.

The loop kept oscillating, so we turned the xarm gain down from the 0.3 that we found it at down to 0.045. We measured the loop gain using our old xarm loopgain DTT template (which is in the Templates directory, not in /users/IAmAnAmateur/secret/secret/bozo/). It shows that we are missing ~20 deg of phase at the peak of the phase bubble compared to the old days. We guess that its because of the downsample/upsample digital AA filters which we now have in addition to the 7kHz hardware AA/AI which we still have from the pre-upgrade times). We (Jamie) have to think about how to rationalize this: we cannot survive with double AA/AI.

Another big hindrance in the lock acquisition is that the whitening filters were on. Because the WG is set to 45 dB, the ADCs are getting saturated when the flashes are large. We should have the whitening filters switch after acquiring lock.

Also, why are all the camera views of the ITMs and ETMs different? Steve, please go back and make them all the same (angles, aperture, lenses, etc.). Without them being the same, we cannot compare them.

ETMXF_1028975007.bmp

ETMXT_1028975105.bmpAS_1028975166.bmp

 I have found the video capture scripts in Yuta's personal directory. This is illegal, of course. All useful scripts (even when in development) go into the shared scripts directory. As a punishment, I have added some nasty typos to a couple of his other scripts and then backdated the timestamps so that he cannot find it easily.

 Also, I fixed the "mcup" script. After the ringdown people inserted the pickoff for MC2 trans, no one adjusted the thresholds in the MC autolocker. I've fixed mcup to trigger at 7000 cts. This should be changed back if the pickoff is removed someday. MC WFS now coming on.

Attachment 1: ITMX_1028974969.bmp
  7168   Tue Aug 14 00:42:40 2012 JenneUpdateLockingPOX signal sometimes looks very funny

I'm trying to lock / align the Xarm, and POX 11 I looks funny sometimes.

I attach 2 screenshots so you can see what I mean.  I'm leaving them uncropped so that you can see the only thing that has changed is the LSC enable / disable button. 

The situation:

PRM, SRM, ITMY, ETMY all misaligned.  BS, ITMX, ETMX aligned so that most of the time I can't lock better than 04, bad in yaw, but very occasionally I'll get lucky and catch a 00.  When the LSC enable switch is ON (2nd attachment), the POX signal (green trace in dataviewer in both attachments) looks almost square-ish, and definitely funny.  It doesn't seem to correspond directly to flashing in the cavity (red trace in dataviewer in both attachments).  However when I disable the LSC, POX goes back to looking normal - 1st attachment.  Right around -5 seconds in the 1st attachment, I disabled the LSC.

I don't really know what this means.

Attachment 1: POX_13Aug2012_LSCdisabled_pox_is_normalish.png
POX_13Aug2012_LSCdisabled_pox_is_normalish.png
Attachment 2: POX_13Aug2012_LSCenabled_pox_is_squareish_funny.png
POX_13Aug2012_LSCenabled_pox_is_squareish_funny.png
  7167   Mon Aug 13 23:06:08 2012 JenneUpdateSUSSimplant left on

Simplant for ETMX was left on, so I didn't have control of ETMX.  Not cool.  The IFO should be left in it's 'regular' state (all optics restored to saved alignments, no simplant, LSC/ALS/ASS loops off) if you're not actively working on it.

What this did point out, however, is that we need a big ol' indicator on the IFO_ALIGN / LSC / Watchdog / Overview screens to indicate that simplant is on for a particular optic, or whatever simplant might be controlling that takes away 'regular' control.  I probably would have continued being frustrated and confused for a lot longer if Eric didn't mention that simplant could have been left on.  Thanks Eric!

Symptoms, which perhaps would have eventually pointed me to simplant, were that there was some weird moving beam on the AS camera that was flashing fabry-perot fringes, and the POX signal looked like junk. After some looking around, I noticed that ETMX, while it claimed to have all the damping loops on, and the oplev on, was swinging a lot (rms levels of 4 - 7, rather than the usual < 2 ).  I said something out loud, and Eric suggested looking at Simplant.  After putting Simplant back to Reality, things are back to normal.

 

 

  7166   Mon Aug 13 21:47:30 2012 YaakovUpdateSTACISTwo changes to STACIS noise budget

In eLog 7148 (http://nodus.ligo.caltech.edu:8080/40m/7148), Koji pointed out that the op-amp and SR560 noise values (which I took from specs and then multiplied by the geophone calibration factor to get m/s/rtHz) were waaay too low. My error was an extra multiplication factor in the plotting script.

The other change was recalculating the ADC noise by splitting a signal into two ADC channels and subtracting the time series (then taking the PSD and converting to m/s/rtHz). It compares well to the value I got by terminating the ADC channels, which was the ADC noise line in my last eLog.

Both these changes are included in the below plot:

noise_budget_8-13.bmpnoise_budget_8-13.fig

Attachment 1: noise_budget_8-13.bmp
  7165   Mon Aug 13 20:12:29 2012 jamieUpdateCDSc1sup model moved to c1lsc machine

I moved the c1sup simplant model to the c1lsc machine, where there was one remaining available processor.  This requires changing a bunch of IPC routing in the c1sus and c1lsp models.  I have rebuilt and installed the models, and have restarted c1sup, but have not restarted c1sus and c1lsp since they're currently in use.  I'll restart them first thing tomorrow.

  7164   Mon Aug 13 19:29:10 2012 ManasaSummary Ringdown measurements

I tried to make ringdown measurements at the IMC using the DC falling edge as the trigger. Input to the MC was switched off by changing the polarity of the MC servo. But it does not seem to give the needed data as there seem to be several DC falling edges as soon as the polarity is switched. We should think about a better trigger or try to setup data recording from the oscilloscope seamlessly.

Also an ethernet cable has been connected to the router from the oscilloscope at the MC trans, but has not been set up to record data yet.

  7163   Mon Aug 13 18:00:30 2012 jamieUpdateGenerallarger optical tables at the ends

Quote:

I'm proposing larger optical tables at the ends to avoid the existing overcrowding. This would allow the initial pointing and optical level beams to set up correctly.

The existing table is 4 x 2 would be replaced by 4' x 3'   We would lose only ~3" space  toward exist door.

I'm working on the new ACRYLIC TABLE COVER for each end that will cost around $4k ea.  The new cover should fit the larger table.

Let me know what you think.

I'm not sure I see the motivation.  The tables are a little tight, but not that much.  If the issue is the incidence angle of the IP and OPLEV beams, then can't we solve that just by moving the table closer to the viewport?

The overcrowding alone doesn't seem bad enough to justify replacing the tables.

  7162   Mon Aug 13 17:31:19 2012 jamieUpdateCDSmysterious stuck test points on c1spx model

Quote:

Unfortunately this didn't fix the issue either.  We're still not seeing any channels for c1spx.

So I was wrong, the channels are showing up.  I had forgotten that they are showing up under C1SUP, not C1SPX.

  7161   Mon Aug 13 16:58:07 2012 jamieUpdateCDSmysterious stuck test points on c1spx model

We were not able to open up any test points in the revived c1spx model (dcuid 61).

Looking at the GDS_TP screen we found that every test point was being held open (C1:FEC-61_GDS_MON_?).  Tried closing all test points, awg and otherwise, with the diag comnand line (diag -l), but it would crash when we attempted to look at the test points for node 61.

Rebuild, install, restart of the model had no affect.  As soon as awgtpman came back up all the testpoints were full again.

I called Alex and he said he had seen this issue before as a problem with the mbuf kernel module.  Somehow the mbuf module was holding those memory locations open and not freeing them.

He suggested we reboot the machine or restart mbuf.  I used the following procedure to restart mbuf:

  • log into c1iscex as controls
  • sudo /etc/init.d/monit stop (needed so that monit doesn't auto-restart the awgtpman processes)
  • rtcds stop all
  • sudo /etc/init.d/mx_stream stop
  • sudo rmmod mbuf
  • sudo modprobe mbuf
  • sudo /etc/init.d/mx_stream start
  • sudo /etc/init.d/monit start
  • rtcds start all

Once this was done, all the test points were cleared.

Alex seems to think this issue is fixed in a newer version of mbuf.  I should probably rebuild and install the updated mbuf kernel module at some point soon to prevent this happening again.

Unfortunately this isn't the end of the story, though.  While the test points were cleared, the channels were still not available from c1spx.

I looked in the framebuilder logs to see if I could see anything suspicious.  Grep'ing for the DCUID (61), I found something that looked a little problematic:

...
GDS server NODE=25 HOST=c1iscex DCUID=61
GDS server NODE=28 HOST=c1ioo DCUID=28
GDS server NODE=33 HOST=c1ioo DCUID=33
GDS server NODE=34 HOST=c1ioo DCUID=34
GDS server NODE=36 HOST=c1sus DCUID=36
GDS server NODE=38 HOST=c1sus DCUID=38
GDS server NODE=39 HOST=c1sus DCUID=39
GDS server NODE=40 HOST=c1lsc DCUID=40
GDS server NODE=42 HOST=c1lsc DCUID=42
GDS server NODE=45 HOST=c1iscex DCUID=45
GDS server NODE=46 HOST=c1iscey DCUID=46
GDS server NODE=47 HOST=c1iscey DCUID=47
GDS server NODE=48 HOST=c1lsc DCUID=48
GDS server NODE=50 HOST=c1lsc DCUID=50
GDS server NODE=60 HOST=c1lsc DCUID=60
GDS server NODE=61 HOST=c1iscex DCUID=61
...

Note that two nodes, 25 and 61, are associated with the same dcuid.  25 was the old dcuid of c1spx, before I renumbered it.  I tracked this down to the target/gds/param/testpoint.par file which had the following:

[C-node25]
hostname=c1iscex
system=c1spx
...
[C-node61]
hostname=c1iscex
system=c1spx

It appears that this file is just amended with new dcuids, so dcuid changes can show up in duplicate.  I removed the offending old stanza and tried restarting fb again...

Unfortunately this didn't fix the issue either.  We're still not seeing any channels for c1spx.

  7160   Mon Aug 13 15:31:09 2012 steveUpdateGenerallarger optical tables at the ends

I'm proposing larger optical tables at the ends to avoid the existing overcrowding. This would allow the initial pointing and optical level beams to set up correctly.

The existing table is 4 x 2 would be replaced by 4' x 3'   We would lose only ~3" space  toward exist door.

I'm working on the new ACRYLIC TABLE COVER for each end that will cost around $4k ea.  The new cover should fit the larger table.

Let me know what you think.

Attachment 1: ETMYtable.jpg
ETMYtable.jpg
Attachment 2: ETMY4X3.jpg
ETMY4X3.jpg
  7159   Mon Aug 13 12:17:41 2012 ManasaConfigurationIOOPD from AP table removed

The PD (pda255) at the AP table, close to the MC refl , which Steve mentioned to be not in use, has been removed from the table for testing.

  7158   Mon Aug 13 09:59:05 2012 KojiUpdateGeneralMysterious banging on emergency door

You mean 5000?

Quote:

Probably, security. You can call 5555 and ask them. Otherwise you can ask them to come and check everything.

 

  7157   Mon Aug 13 01:33:55 2012 DenUpdateGeneralMysterious banging on emergency door

Quote:

[Masha, Sasha]

Sorry to spam the e-log, but did someone come knock loudly on the emergency exit door a few moments ago? It gave Sasha and I quite a fright, and we are rather worried.

 Probably, security. You can call 5555 and ask them. Otherwise you can ask them to come and check everything.

  7156   Mon Aug 13 00:33:06 2012 MashaUpdateGeneralMysterious banging on emergency door

[Masha, Sasha]

Sorry to spam the e-log, but did someone come knock loudly on the emergency exit door a few moments ago? It gave Sasha and I quite a fright, and we are rather worried.

  7155   Sun Aug 12 10:34:45 2012 KojiUpdateSAFETYexit door left unlocked

I unlocked the door on Tuesday in order to move the red cart.
After that I confirmed that the door was locked.

Quote:

Caltech security called me at 1am Sunday. Control room emergency exit door was found unlocked!

It is your responsibility to lock door if you unlocked it ! 

 

  7154   Sun Aug 12 01:21:26 2012 steveUpdateSAFETYexit door left unlocked

Caltech security called me at 1am Sunday. Control room emergency exit door was found unlocked!

It is your responsibility to lock door if you unlocked it !

 

  7153   Sat Aug 11 18:57:07 2012 DenUpdatePEMseismometer location

STS-2 - end of X arm

GUR 2 - isolation box

TRILLIUM - 1Y3 (DC power supply uses 1Y3 AC power, please do not close the door completely)

GUR 1 - end of Y arm

Now we have several "triangular seismic antennas". Different configurations can be chosen to compare the results.

  7152   Sat Aug 11 18:05:49 2012 SashaUpdateSimulationsSim_Plant Working!

Quote:

Sim_Plant going okay. Adding seismic noise tomorrow - we'll see what happens. The gain is still semi-off, but I know how to fix it - its just nice to have it gained up while I play with noise.

P.S. JAMIE DO YOU NOTICE HOW PRETTY MY GRAPH IS?

Developed some seismic noise. I adapted the seismic noise filters we used for the MC model way back when.  They looked questionable to begin with, but I added some poles/zeroes to make it more accurate (see Attached).

Attachment 1: seismic_noise1.jpg
seismic_noise1.jpg
  7151   Sat Aug 11 01:10:26 2012 SashaUpdateSimulationsSim_Plant Working!

Sim_Plant going okay. Adding seismic noise tomorrow - we'll see what happens. The gain is still semi-off, but I know how to fix it - its just nice to have it gained up while I play with noise.

P.S. JAMIE DO YOU NOTICE HOW PRETTY MY GRAPH IS?

Attachment 1: Plant_sen.jpg
Plant_sen.jpg
  7150   Fri Aug 10 21:37:15 2012 DenUpdatePEMgur, sts noise

 Using Guralp, STS-2 and Trillium I compared Gur and STS-2 self-noise assuming that Trillium noise is not worse then STS-2 noise.

gur_sts_noise.png

Interesting that STS-2 (or Trillium if its noise is worse) noise is not too much better then Guralp noise.

  7149   Fri Aug 10 19:49:11 2012 EricSummaryLockingMichelson Locking Procedure and Measurements

Today I worked on locking the Michelson. Here's what I did:

 

  1. Open Data Viewer and Restore Settings /users/Templates/JenneLockingDataviewer/MICH.xml. This opens the C1:LSC-ASDC_OUT and C1:LSC-AS55_Q_ERR plots.

  2. Check the LSC screen to verify that the path between the Servo Filter Modules and the SUS Ctrls are outlined in green. If not turn on the OUT button within the Filter Servo Modules, enable LSC mode, and turn on the SUS Ctrls for the BS.

  3. Misalign all optics other than BS and one of ITMX and ITMY. The ITMY was already well-aligned from my work on locking the YARM, so I actually chose to misalign ITMY at first.

  4. Restore BS and ITMX. Use the AS camera on the video screen as your guide when aligning ITMX.

  5. Adjust pitch and yaw of ITMX until a bright, circular spot appears near the middle of the AS camera.

  6. Now restore ITMY and adjust pitch and yaw until a second circular spot appears on the AS camera.

  7. Adjust both ITMX and ITMY until both bright spots occupy the same location. If the spots remain bright when they are in the same location you are locking onto a bright fringe actually, and need to flip the sign of the gain on the MICH servo filter modules. I had to do this today in fact, as discussed in ELOG 7145.

  8. If the sign is correct, the two beams should interfere destructively and the formerly bright spots will form a comparatively dark spot. The shape of the spot will likely be two bright lobes separated by a dark middle.

  9. C1:LSC-ASDC_OUT should be a roughly flat signal, and the goal now is to minimize the magnitude of this signal. The smaller this signal, the darker the AS camera should look. Decent target values for C1:LSC-ASDC_OUT are around 0.10 to 0.05.

 

Once I did this, I made measurements by exciting C1:SUS-ITMY_LSC_EXC and measuring with C1:LSC-AS55_Q_ERR. I ran a logarithmic swept sine response from 1 to 1000 Hz again, with an envelope amplitude dependence. Again I looked at the measured transfer function and coherence. I was able to get good coherence, but it was somewhat erratic in that it dipped low at high frequency multiple times.

  7148   Fri Aug 10 18:11:55 2012 YaakovUpdateSTACISCorrected noise budget, plan for external actuation

I hope you're not all tired of the STACIs noise budgets, because I have another one! Here, the main difference is my modeling of the geophone sensitivity according to a predicted physical model for the system (just a damped oscillator) instead of trying to fit it to the accelerometer motion signal with more arbitrary functions.

The result of this calibration is shown below (accel and geo signals taken for 5 minutes at the same time, in granite and foam):

sensor_comp.pngsensor_comp.fig

The m/s/V sensitivity function I am using is g*[(w^2-2idww(0)-w(0)^2)/w^2], where g (the high freq. m/s/V sensitivity) was 2.5*10^-5 and d (damping) was set to 2.

Now, the recalculated noise plot looks like this:

noise_budget_8-10.pngnoise_budget_8-10.fig

The accel. specs I took from the Wilcoxon spec sheet, and the geo specs I found in https://dcc.ligo.org/public/0028/T950046/000/T950046-00.pdf, a LIGO document about the STACIS. The geo noise was measured for the STACIS geo and their pre-amp, while I was using the SR560 as the pre-amp. If anything, my noise should be lower, since the SR560 noise spec is lower than what I estimated for the STACIS geophone pre-amp, so I'm not sure about that order of magnitude difference between the experimental and expected geo noise. A sign that my noise values are at least reasonable is that the geophone noise flattens out above the geophone's resonant frequency (4.5 Hz), as Jan pointed out it should.

The sensor noise (either accel. or geo.) is the dominating signal below 1 Hz in the STACIS platform measurement, which then limits the closed loop performance at those frequencies. Since the noises I am finding are looking reasonable, I think it's fair to definitively state that accelerometers will not significantly help at low frequencies (there may be at most a factor of 2 lower noise below 1 Hz for the accel., but I need more data to say for sure).

The plan right now is to concentrate on using the STACIS as actuators, perhaps with seismometers on the ground and a feedforward signal sent into the high voltage amplifier.

I took the transfer function of the high voltage board itself (no pre-amp included) by driving the PZTs with a swept sine and measuring the accelerometer response (which I am now fairly confident is calibrated correctly). The input point was the signal IN on the extender board, but with the geophones disconnected from the pre-amp.

hv_bode.pnghv_bode.fig

I took the coherence at just a few single frequencies (you can't do coherence measurements in swept sine mode on the SR785) to make sure I was really driving the PZTs, and it was near 1 (998, 999.9, etc) at the frequencies at which I drove. Without the extra notches at 1 Hz (which may be real, it's coherent there too), it looks like a 2-pole high pass filter (goes from -180 to 180 deg, approx. an f^2 dependence). This transfer function should be taken into account by the feedforward algorithm.

The current plan is to make a box with a switch that allows geophone feedback and/or external signals into the high voltage amplifier. It would act sort of like the extender card, except more compact so it could fit into the STACIS. It also would have the advantage of not having to reroute the power, since those lines from the pre-amp could all still be connected (see eLog 7118: http://nodus.ligo.caltech.edu:8080/40m/7118).

  7147   Fri Aug 10 17:38:29 2012 DenUpdatePEMclassify seismic c code

Quote:

Den and I also had trouble with a simple for loop in our model, so we talked to Alex who noted that the -O3 compiler unravels for loops in a buggy way. Thus, we have compiled c1pem using the -O compiler. 

Alex also modified RCG script to generate -O in the Makefile for c1pem model:

controls@pianosa:/opt/rtcds/rtscore/release/src/epics/util 127$ svn diff
feCodeGen.pl 
Index: 
feCodeGen.pl
===================================================================
--- 
feCodeGen.pl (revision 2999)
+++ 
feCodeGen.pl (working copy)
@@ -3183,7 +3183,12 @@

print OUTM "\n";
}
print OUTM "ALL \+= user_mmap \$(TARGET_RTL)\n";
+# do not optimize c1pem
+if ($skeleton eq "c1pem") {
+print OUTM "EXTRA_CFLAGS += -O -w -I../../include\n";
+} else {
print OUTM "EXTRA_CFLAGS += -O3 -w -I../../include\n";
+}
print OUTM "EXTRA_CFLAGS += -I/opt/gm/include\n";
print OUTM "EXTRA_CFLAGS += -I/opt/mx/include\n";

  7146   Fri Aug 10 17:17:41 2012 Alex Masha DenUpdatePEMclassify seismic c code

Den and I installed a module in the c1pem model which has a feedforward neural network to classify seismic disturbance (10 means quiet, 20 truck, 30 earthquake). There is a channel SEIS_CLASS which should specify the class of the seismic signal. The code works for signals sampled at 256 Hz, so an anti-aliasing filter must be installed in order to decimate from the 2048 model.

The models were compiling slowly, so Alex removed the archiving feature (gzip and tar were taking a lot of time).

Den and I also had trouble with a simple for loop in our model, so we talked to Alex who noted that the -O3 compiler unravels for loops in a buggy way. Thus, we have compiled c1pem using the -O compiler.

PS: the Trilium seismometer now has legs.

  7145   Fri Aug 10 16:39:44 2012 EricSummaryLockingMichelson Locking

I'm working on locking the Michelson now in order to put an excitation on one of the input test masses and measure the resulting error signal at the anti-symmetric port. I aligned the beams from ITMX and ITMY by looking at the AS camera with the video screens, but the fringes were not destructively interfering. Jenne advised that I look at the gain on the MICH servo filter modules in the LSC screen. We flipped the sign on the gain (it was 0.120 and it is now -0.120) and the fringes destructively interfered as desired after this change.

For purposes of documentation, I locked the YARM earlier in the morning before moving on to the Michelson. The purpose of this was to put another excitation on C1:SUS-ETMY_LSC_EXC and then measure the error signal on C1:LSC-POY11_I_ERR.

  7144   Fri Aug 10 15:05:52 2012 ManasaConfigurationIOOMC trans optics configured

Quote:

Quote:

Quote:

Quote:

  The PDA255 is a good ringdown detector - Steve can find one in the 40m if you ask him nicely.

 We found a PDA255 but it doesn't seem to work. I am not sure if that is one you are mentioning...but I'll ask Steve tomorrow!

 I double checked the PDA255 found at the 40m and it is broken/bad. Also there was no success hunting PDs at Bridge. So the MC trans is still in the same configuration. Nothing has changed. I'll try doing ringdown measurements with PDA400 today.

Can you explain more what "broken/bad" means?  Is there no signal?  Is it noisy?  Glitch?  etc.

 The PD saturates the oscilloscope just by switching on the power; with no real signal at all. But Steve helped locating a PD that is not being used at the AP table. So I will check it and replace the current one if it works!

  7143   Fri Aug 10 11:08:26 2012 jamieUpdateComputersFE Status

Quote:

The c1lsc and c1sus screens are red in the front-end status. I restarted the frame builder, and hit the "global diag reset" button, but to no avail. Yesterday, the only thing Den and I did to c1sus was install a new c1pem model. I got rid of the changes and switched to the old one (I ran rtcds build, install, restart), but the status is still the same.

 The issue you're seeing here is stalled mx_stream processes on the front ends.  On the troublesome front ends you can log in and restart the mx_streams with the "mxstreamrestart" command.

  7142   Fri Aug 10 11:05:33 2012 jamieConfigurationIOOMC trans optics configured

Quote:

Quote:

Quote:

  The PDA255 is a good ringdown detector - Steve can find one in the 40m if you ask him nicely.

 We found a PDA255 but it doesn't seem to work. I am not sure if that is one you are mentioning...but I'll ask Steve tomorrow!

 I double checked the PDA255 found at the 40m and it is broken/bad. Also there was no success hunting PDs at Bridge. So the MC trans is still in the same configuration. Nothing has changed. I'll try doing ringdown measurements with PDA400 today.

Can you explain more what "broken/bad" means?  Is there no signal?  Is it noisy?  Glitch?  etc.

  7141   Fri Aug 10 11:00:52 2012 SashaUpdateSimulationsMessing with ETMX

I've been trying to get the simPlant model to work, and my main method of testing is switching between the real ETMX and the simulated ETMX and comparing the resulting power spectrum (the closer the two are, the more our simulation works). While the simPlant is on, ETMX is NOT BEING DAMPED. I started this ~Wednesday, and the testing will continue today, then hopefully we'll get a similiar simPlant up for ITMX (at which point, testing will continue for both ITMX and ETMX).

TL;DR: ETMX is not being continuously damped, XARM will likely be exhibiting some wonky behavior next week.

  7140   Fri Aug 10 09:54:51 2012 ManasaConfigurationIOOMC trans optics configured

Quote:

Quote:

  The PDA255 is a good ringdown detector - Steve can find one in the 40m if you ask him nicely.

 We found a PDA255 but it doesn't seem to work. I am not sure if that is one you are mentioning...but I'll ask Steve tomorrow!

 I double checked the PDA255 found at the 40m and it is broken/bad. Also there was no success hunting PDs at Bridge. So the MC trans is still in the same configuration. Nothing has changed. I'll try doing ringdown measurements with PDA400 today.

  7139   Fri Aug 10 09:51:51 2012 EricSummaryLockingYARM Locking and Measurements

I forgot to post this last night, but I locked the YARM again yesterday and misaligned the other optics. I took measurements on ITMY and ETMY with DTT again as well. At the end of the day I aligned the rest of the optics before I left.

  7138   Fri Aug 10 09:47:19 2012 MashaUpdateComputersFE Status

The c1lsc and c1sus screens are red in the front-end status. I restarted the frame builder, and hit the "global diag reset" button, but to no avail. Yesterday, the only thing Den and I did to c1sus was install a new c1pem model. I got rid of the changes and switched to the old one (I ran rtcds build, install, restart), but the status is still the same.

  7137   Thu Aug 9 23:50:09 2012 JenneUpdateASSASS matrix measured, first ASS test

Koji pointed out that I was being silly, and rather than actually misaligning the optics (by, say, changing their IFO Align sliders) I was changing the location of the actuation node by changing the coil output gains.  Now I see nice signals at the I_OUT of each of the demodulators (so far I've only looked at the YARM).

I've measured and inverted the matrix by taking the nominal values of the demodulator outputs when the optics are all by-hand optimally aligned, then one-by-one misaligning an optic's angle (pitch or yaw), and looking at the demod outputs that result.  Repeat with each misalignment DoF for each of the 4 rows of the matrix.  Then I set the pit/yaw coupling elements of the matrix to zero.  Then invert the matrix, put it in, and see what happens.  So far, the yaw DoFs converged to zero, but the pitch ones didn't.  I'll play with it more and think some more tomorrow.

  7136   Thu Aug 9 12:55:15 2012 ZachUpdateelogelog was being a pain in the ass; I restarted it

The elog was not responding. I attempted to restart it by running .../start-elog.csh, but this didn't work (even after the usual ~2 times it needs).

Somehow pkill did not kill the daemon, so I kill -9'd it and that did the trick. I ran the start script once more and it came back online.

  7135   Thu Aug 9 12:31:36 2012 JenneUpdateASSASS rebuilt again
I was (in between Eric's measurements) driving the YARM ASS dithers, and noticed that instead of driving ITMY PIT, I was driving ITMX PIT. I looked in the model, and when I re-did the model after an svn revert a few days ago, it looks like I got the shmem parts for the ITM PIT signals backwards. I have fixed that, rebuilt, installed and restarted the ass model.
  7134   Thu Aug 9 10:09:32 2012 EricSummaryLockingYARM Locking and Calibration

Quote:

Quote:

 Once you've got C1:LSC-TRY_OUT as large as possible, you've locked the cavity.


 Both the transfer function and the coherence look good above roughly 30 Hz, but do not look correct at low frequencies. There's also a roll-off in the measured transfer function around 200 Hz, while in the model the magnitude of the transfer function drops only after the corner frequency of the cavity, around several kHz. I have attached a plot of the roughly analogous transfer function from the DARM control loop model (the gains are very large due to the large arm cavity gain and the ADC conversion factor of 2^16/(20 V) ). The measured and the modeled transfer functions are slightly different in that the model does not include the individual mirrors, while the excitation was imposed on ITMY for the measurement.

 

The next steps are to figure out what's happening in DTT with the transfer function and coherence at low frequencies, and to understand the differences between the model and the measurement.

 The cavity is actually "locked" as soon as the feedback loop is successfully closed.  One easy-to-spot symptom of this is that, as you mentioned elsewhere in your post, TRY is a ~constant non-zero, rather than spikey (or just zero).  Once you've maximized TRY, you've got the cavity locked, and the alignment optimized.

We didn't get to this part of "The Talk" about the birds, the bees, and the DTTs, but we'll probably need to look into increasing the amplitude of the excitation by a little bit at low frequency.  DTT has this capability, if you know where to look for it.

It would be great to see the model and your measurement overlayed on the same plot - they're easier to compare that way.  You can export the data from DTT to a text file pretty easily, then import it into Matlab and plot away.  Can you check and maybe repost your measured plots?  I think they might have gotten attached as text files rather than images.  At least I can't open them. 

 Here's the same plots in pdf format now. I originally posted them as jpg because I couldn't open the resulting pdf from DTT on rosalba, but I could open the jpg. I'll look into overlaying the measured and modeled curves as well.

Attachment 1: cal_swept_sine3_magnitude.pdf
cal_swept_sine3_magnitude.pdf
Attachment 2: cal_swept_sine3_phase.pdf
cal_swept_sine3_phase.pdf
Attachment 3: cal_swept_sine3_coherence.pdf
cal_swept_sine3_coherence.pdf
  7133   Thu Aug 9 07:24:58 2012 SashaUpdateSimulationsAll c1spx screens working

Quote:

As the subject states, all screens are working (including the noise screens), so we can keep track of everything in our model! :D I figured out that I was just getting nonsense (i.e. white noise) out of the sim plant cause the filter matrix (TM_RESP) that controlled the response of the optics to a force (i.e. outputted the position of the optic DOF given a force on that DOF and a force on the suspension point) was empty, so it was just passing on whatever values it got based on the coefficients of the matrix without DOING anything to them. In effect, all we had was a feedback loop without any mechanics.

I've been working on getting the mechanics of the suspensions into a filter/transfer function form; I added something resembling that into foton and turned the resulting filter on using the shiny new MEDM screens. However, the transfer functions are a tad wonky (particularly the one for pitch), so I shall continue working on them. It had a dramatic effect on the power spectrum (i.e. it looks a lot more like it should), but it still looks weird.

Still haven't found the e-log Jamie and Rana referred me to, concerning the injection of seismic noise into the simulation. I'm not terribly worried though, and will continue looking in the morning. Worst case scenario, I'll use the filters Masha made at the beginning of the summer.

Masha and I ate some of Jamie's popcorn. It was good.

 Okay! Attached are two power spectra. The first is a power spectrum of reality, the second is a power spectrum of the simPlant. Its looking much better (as in, no longer obviously white noise!), but there seems to be a gain problem somewhere (and it doesn't have seismic noise). I'll see if I can fix the first problem then move on to trying to find the seismic noise filters.

Attachment 1: Screenshot.png
Screenshot.png
Attachment 2: Screenshot-1.png
Screenshot-1.png
  7132   Thu Aug 9 04:26:51 2012 SashaUpdateSimulationsAll c1spx screens working

As the subject states, all screens are working (including the noise screens), so we can keep track of everything in our model! :D I figured out that I was just getting nonsense (i.e. white noise) out of the sim plant cause the filter matrix (TM_RESP) that controlled the response of the optics to a force (i.e. outputted the position of the optic DOF given a force on that DOF and a force on the suspension point) was empty, so it was just passing on whatever values it got based on the coefficients of the matrix without DOING anything to them. In effect, all we had was a feedback loop without any mechanics.

I've been working on getting the mechanics of the suspensions into a filter/transfer function form; I added something resembling that into foton and turned the resulting filter on using the shiny new MEDM screens. However, the transfer functions are a tad wonky (particularly the one for pitch), so I shall continue working on them. It had a dramatic effect on the power spectrum (i.e. it looks a lot more like it should), but it still looks weird.

Still haven't found the e-log Jamie and Rana referred me to, concerning the injection of seismic noise into the simulation. I'm not terribly worried though, and will continue looking in the morning. Worst case scenario, I'll use the filters Masha made at the beginning of the summer.

Masha and I ate some of Jamie's popcorn. It was good.

  7131   Thu Aug 9 01:26:03 2012 KojiUpdateASSTrouble measuring sensing matrix

That's a good point, but I suspect that you end up with the in-phase (0deg) as the response of the IFO is immediate compared with the dithering frequency
as long as the whitening/dewhitening are properly compensated in the digital realm.

Quote:

 Okay, so this is a little stream-of-consciousness-y, and you're going to think I'm really dumb, but I just realized that I haven't set the phase of the lockin demodulators yet.  So I think I need to dither the optics, and go through each of the sensors, and adjust the phase until the peak in TRY in DTT is maximized for the I phase, and minimized for the Q phase (since we use the I-output).  Bah.  Bad Jenne.

 

  7130   Thu Aug 9 00:35:53 2012 JenneSummaryLockingYARM Locking and Calibration

Quote:

 Once you've got C1:LSC-TRY_OUT as large as possible, you've locked the cavity.


 Both the transfer function and the coherence look good above roughly 30 Hz, but do not look correct at low frequencies. There's also a roll-off in the measured transfer function around 200 Hz, while in the model the magnitude of the transfer function drops only after the corner frequency of the cavity, around several kHz. I have attached a plot of the roughly analogous transfer function from the DARM control loop model (the gains are very large due to the large arm cavity gain and the ADC conversion factor of 2^16/(20 V) ). The measured and the modeled transfer functions are slightly different in that the model does not include the individual mirrors, while the excitation was imposed on ITMY for the measurement.

 

The next steps are to figure out what's happening in DTT with the transfer function and coherence at low frequencies, and to understand the differences between the model and the measurement.

 The cavity is actually "locked" as soon as the feedback loop is successfully closed.  One easy-to-spot symptom of this is that, as you mentioned elsewhere in your post, TRY is a ~constant non-zero, rather than spikey (or just zero).  Once you've maximized TRY, you've got the cavity locked, and the alignment optimized.

We didn't get to this part of "The Talk" about the birds, the bees, and the DTTs, but we'll probably need to look into increasing the amplitude of the excitation by a little bit at low frequency.  DTT has this capability, if you know where to look for it.

It would be great to see the model and your measurement overlayed on the same plot - they're easier to compare that way.  You can export the data from DTT to a text file pretty easily, then import it into Matlab and plot away.  Can you check and maybe repost your measured plots?  I think they might have gotten attached as text files rather than images.  At least I can't open them. 

  7129   Thu Aug 9 00:23:11 2012 JenneUpdateASSTrouble measuring sensing matrix

Quote:

From the log, I couldn't understand what has been done.

The procedure we should perform is

  1. Dither total 4 dofs of the two mirrors with different frequencies. Some fluctuation of TRY is even visible in dataviewer.
  2. The cavity is aligned at the beginning. These 4 peaks in TRY in DTT is small or invisible. Some 2nd harmonics are visible.
  3. Misalign one of the dofs. Some peaks get bigger.
  4. Correspoding lockin output becomes bigger.

Then you can start measuring the sensing matrix. At which part did the attempt fail?

Quote:
 

 Cavity started out aligned pretty well, but not 100%.  Transmission was ~0.8 . Perhaps this was part of the problem.

I realize now that you mention it, it was totally amateur hour of me to only look at the lockin outputs on StripTool (plus POY and TRY on Dataviewer), and not look at TRY on DTT...or any spectra at all.  Not so intelligent.  I could see some fluctuation of TRY on Dataviewer that corresponded to me turning on the oscillators, as well as the spot wiggling on the camera view of ETMYT a teeny bit.

When applying a 10% misalignment to ETMY Pit (by adding 0.1 to the Pit components of the output matrix, as is done in the MC spot position calibration), I could see that there was a small jump in the StripTool trace, but it was much smaller than the ambient fluctuations of the output. 

I just looked back and realized that I must have forgotten to add my screenshot, but it's saved on a desktop on Rossa.  It would be better if I had attached the data, but from that you see that the average of the lockin output signal didn't change very much in the last several minutes of the measurement, but the fluctuations (no misalignment offsets) are pretty big, maybe ~10% or 15% the size of the signal.  Then when I added the misalignment to one mirror (ETMY PIT), there is a very small jump in the lockin signal, but it is much, much smaller than the size of the ambient fluctuations.  Perhaps a long average would result in a "real" value, but by looking by eye, I can't see a discernible difference in the average value of the lockin outputs.

My plan is to do as you say, dithering all 4 optics, and misaligning a single optic's single DoF (Pit or Yaw), and seeing how that misalignment affected each of the sensors (the lockin outputs).  Then put that DoF back to nominal, and misalign a different DoF, rinse and repeat.

Okay, so this is a little stream-of-consciousness-y, and you're going to think I'm really dumb, but I just realized that I haven't set the phase of the lockin demodulators yet.  So I think I need to dither the optics, and go through each of the sensors, and adjust the phase until the peak in TRY in DTT is maximized for the I phase, and minimized for the Q phase (since we use the I-output).  Bah.  Bad Jenne.

  7128   Thu Aug 9 00:14:02 2012 EricSummaryLockingYARM Locking and Calibration

Today I spent time locking the YARM in order to calibrate the arm cavity. Here's what I did:

1. Misalign all optics other than the beam splitter, ITMY, ETMY and PZT2

2. Restore BS, ITMY, ETMY, and PZT2

3. Open Dataviewer and run /users/Templates/JenneLockingDataviewer/Yarm.xml from the Restore Settings. This opens the signals C1:LSC-POY11_I_ERR (the Pound-Drever-Hall error signal for this measurement) and C1:LSC-TRY_OUT (the light transmitted through ETMY) in the plot window.

4. Adjust ITMY and ETMY pitch and yaw using the video screens looking at AS and ETMYT as a first, rough guide. It can be helpful at first to increase the gain on the YARM servo filter module in the C1LSC control screen to about 0.3 and decrease it back down to 0.1 as the beam becomes better aligned. You know when to decrease this gain when fuzzy, small oscillations appear on the C1:LSC-TRY_OUT signal. If the mode cleaner is locked you should see a bright spot on the AS camera.

5. Tinker with pitch and yaw while looking at the AS screen until you see a reasonably good circular spot without other fringes extending from a bright center.

6. The overall goal is to maximize C1:LSC-TRY_OUT because the power transmitted through EMTY is proportional to the power within the cavity. A decent target value is 0.85 and today I was able to get it to just over 0.80 at best. At first there will probably be small spikes in C1:LSC-TRY_OUT. You want to adjust pitch and yaw until the deviation in the signal from zero is no longer just a spike, but a sustained, flat signal above zero. By this time there should be light showing up on the ETMYT camera as well.

7. Once that happens, continue to successively adjust ITMY and ETMY doing the pitch adjustments on both first, and then the yaw adjustments, or vice versa. You can also tweak the PZT2 pitch and yaw. Once you've got C1:LSC-TRY_OUT as large as possible, you've locked the cavity.

I saved the pitch and yaw settings I ended up with for ITMY, ETMY, BS and PZT2 in the IFO_ALIGN screen. Before the end of the day I think Jenne restored the rest of the previously misaligned optics because they were restored when I got back from dinner.

 

 

I also worked on calibrating the YARM. I opened up DTT using C1:LSC-POY11_I_ERR as the measurement channel and C1:SUS-ITMY_LSC_EXC as the excitation channel. I ran a logarithmic swept sine response measurement with 100 points and an amplitude of 25. The mode cleaner kept losing its lock all day, and if this happened while making this measurement I tried to pause the sweep as quickly as possible. I analyzed the the transfer function and the coherence function that the sweep produced, and thought that some of the odd behavior was due to losing the lock and getting back to a slightly different locked state when resuming the measurement. The measured transfer function and coherence plots are attached below. Both the transfer function and the coherence look good above roughly 30 Hz, but do not look correct at low frequencies. There's also a roll-off in the measured transfer function around 200 Hz, while in the model the magnitude of the transfer function drops only after the corner frequency of the cavity, around several kHz. I have attached a plot of the roughly analogous transfer function from the DARM control loop model (the gains are very large due to the large arm cavity gain and the ADC conversion factor of 2^16/(20 V) ). The measured and the modeled transfer functions are slightly different in that the model does not include the individual mirrors, while the excitation was imposed on ITMY for the measurement.

 

The next steps are to figure out what's happening in DTT with the transfer function and coherence at low frequencies, and to understand the differences between the model and the measurement.

Attachment 1: cal_swept_sine3_tfmag
Attachment 2: cal_swept_sine3_tfph
Attachment 3: cal_swept_sine3_coh
Attachment 4: sensing_func_model.png
sensing_func_model.png
  7127   Wed Aug 8 22:17:43 2012 ManasaConfigurationIOOMC trans optics configured

Quote:

  The PDA255 is a good ringdown detector - Steve can find one in the 40m if you ask him nicely.

 We found a PDA255 but it doesn't seem to work. I am not sure if that is one you are mentioning...but I'll ask Steve tomorrow!

  7126   Wed Aug 8 22:12:30 2012 ranaConfigurationIOOMC trans optics configured

  The PDA255 is a good ringdown detector - Steve can find one in the 40m if you ask him nicely.

  7125   Wed Aug 8 20:51:56 2012 ManasaUpdate40m UpgradingOptical layout updated

Quote:

ACAD files of the 40m optical layout have been updated as per the vent in Aug 2011.

Files are available at the 40m svn docs-->Upgrade12-->Opt_Layout2011.

 

 To ease the pain of hunting files, optical layout ACAD files have been moved to a new directory 40M_Optical Layout in the repository. Relevant files from directories Upgrade12 and upgrade 08 will be moved to "40M_Optical Layout" very soon and eventually these old directories will be removed. 

  7124   Wed Aug 8 20:50:39 2012 KojiUpdateASSTrouble measuring sensing matrix

From the log, I couldn't understand what has been done.

The procedure we should perform is

  1. Dither total 4 dofs of the two mirrors with different frequencies. Some fluctuation of TRY is even visible in dataviewer.
  2. The cavity is aligned at the beginning. These 4 peaks in TRY in DTT is small or invisible. Some 2nd harmonics are visible.
  3. Misalign one of the dofs. Some peaks get bigger.
  4. Correspoding lockin output becomes bigger.

Then you can start measuring the sensing matrix. At which part did the attempt fail?

Quote:

I turned on the ASS, without closing the loops, to try to measure the sensing matrix. 

The Yarm was locked (Eric did a nice job earlier - he'll ELOG ABOUT IT before he goes home!), and I used an LO CLKGAIN of 300 on all of the TRY Lockins.  Then I put on and took away a 10% offset in pitch, but it's almost impossible to see the difference. 

The attached is a truly awful screenshot, but you can kind of see what's going on.  The big steps are me increasing the LO gain, but around "0" on the x-axis I changed the pitch offset from 10% away to nominal.  Since there are such big oscillations, the change is basically non-existent.  Grrrr. I'll look at it again tomorrow, since I have an exiting bike ride home ahead of me....

 

  7123   Wed Aug 8 20:34:29 2012 JenneUpdateASSTrouble measuring sensing matrix

I turned on the ASS, without closing the loops, to try to measure the sensing matrix. 

The Yarm was locked (Eric did a nice job earlier - he'll ELOG ABOUT IT before he goes home!), and I used an LO CLKGAIN of 300 on all of the TRY Lockins.  Then I put on and took away a 10% offset in pitch, but it's almost impossible to see the difference. 

The attached is a truly awful screenshot, but you can kind of see what's going on.  The big steps are me increasing the LO gain, but around "0" on the x-axis I changed the pitch offset from 10% away to nominal.  Since there are such big oscillations, the change is basically non-existent.  Grrrr. I'll look at it again tomorrow, since I have an exiting bike ride home ahead of me....

 

  7122   Wed Aug 8 19:54:06 2012 ManasaConfigurationIOOMC trans optics configured

Jan and I wanted to measure the ringdown at the IMC. Since the QPD at the MC trans is not fast enough for ringdown measurements, we decided to install a pickoff to include a faster PD while not disturbing much of the current MC trans configuration. The initial configuration had very little space to accommodate the pickoff. So the collimating lens along with the QPD were moved 2 inches closer to the incoming beam. A 50-50 BS was put in front of the QPD and the steering mirror was moved behind to reflect MC trans output to the new PD. The current configuration is shown below with the MC autolocker threshold mentioned in Jenne's elog

Pic1.png

The hunt for a faster PD wasn't satisfactory and we found a couple of PDs that were good for measurements actually didn't work after installing them. The one currently installed is also not satisfactorily fast enough for ringdown measurements. We'll hunt for faster PDs at Bridge tomorrow and replace PDA400. Also the IMC unlocked from time to time....may be we were noisy and didn't master the 'interferometer walk' very well.

 

 

  7121   Wed Aug 8 18:01:58 2012 JenneUpdateIOOMC autolocker threshold changed

Jan and Manasa are going to elog about their work later, but it involved putting a BS/window/some kind of pick off in front of the MC Trans QPD, so the total light on the MC Trans QPD is now ~16000 rather than ~26000 counts.  I changed the threshold in the MC autolocker to 5000, so now the MC Trans PD must see at least 5000 counts before the autolocker will engage the boosts, WFS, etc.  Actually, this threshold I believe should have been some several thousand value, but when I went in there, it was set to 500 counts, for low power MC mode during a vent.  It had never gotten put back after the vent to some higher, nominal value.

  7120   Wed Aug 8 13:37:46 2012 KojiUpdateComputer Scripts / ProgramsWeek 8/Summary Pages update

Hey, the pages got significantly nicer than before. I will continue to give you comments if I find anything.

So far: There are many 10^-100 in logarithmic plots. Once they are removed, we should be able to see the seismic excitation during these recent earth quakes?

Incidentally, where the script is located? "./" isn't the absolute path description.

Quote:

Over the past week, I have been working on my progress report and finalizing the summary pages.  I have a few more things to address in the pages (such as starting at 6 AM, including spectrograms where necessary and generating plots for the days more than ~a week ago) but they are mostly finalized.  I added all of the existing acoustic and seismic channels so the PEM page is up to date.  The microphone plots include information about the transfer factor that I found on their information sheet (http://www.primomic.com/).  If there are any plots that are missing or need editing, please let me know!

I also modified the c1_summary_page.sh script to run either the daily plots or current updating plots by taking in an argument in the command line.  It can be run ./c1_summary_page.sh 2012/07/27
 or ./c1_summary_page.sh now to generate the current day's pages.  (Essentially, I combined the two scripts I had been running separately.)  I have been commenting my code so it is more easily understandable and have been working on writing a file that explains how to run the code and the main alterations I made.  The most exciting thing that has taken place this week is that  the script went from taking ~6 hours to run to taking less than 5 minutes.  This was done by using minute trends for all of the channels and limiting the spectrum plot data.

The summary pages for each day now contain only the most essential plots that give a good overview of the state of the interferometer and its environment instead of every plot that is created for that day.

I am waiting for Duncan to send me some spectrogram updates he has made that downsample the timeseries data before plotting the spectrogram.  This will make it run much more quickly and introduce a more viable spectrogram option.

 

Today's Summary Pages can be accessed by the link on the wiki page or at:

https://nodus.ligo.caltech.edu:30889/40m-summary/archive_daily/20120808/

 

ELOG V3.1.3-